# Bayes2IMC: In-Memory Computing for Bayesian Binary Neural Networks

### v1

@@ III-A Overview
The parameters $\boldsymbol{p}_{\boldsymbol{w}}=\{p_{w}\}_{w\in\boldsymbol{w}}$ required for sample generation are stored in WP. The noise, provided by the NP, is due to the inherent device-to-device and cycle-to-cycle stochasticity represented by the state-dependent programming noise shown in Fig. 3. This Weight Plane-Noise Plane (WP-NP) crossbar array architecture introduced in [6], assigns $M$ rows for the weight probability parameters, and $L$ rows in the NP to generate the required stochasticity. We reuse the hardware resource for sampling (i.e., $L\ll M$) and empirically show that this does not affect network accuracy, leading to improved area efficiency. Noise cells from NP and weight cells from WP combine to generate $w$. We now elaborate the details of this sampling utilizing device stochasticity.
@@ III-B Leveraging Device Stochasticity
Reparametrization: To generate network parameters as per (9), we require a source of Uniform noise. However, the naturally available stochasticity from the PCM-based NP in the array is Gaussian distributed. One method to utilize the hardware-provided noise is to soft-binarize the natural parameter described in (2) using the Gumbel-Softmax trick [6, 7, 34, 41]. To make this method amenable to implementation on PCM, the stochastic parameter in the soft-binarization term given in [34] would need to be approximated to a Gaussian random variable (RV) by Gaussian fitting, as described in [6]. We develop a more principled method for sampling binary weights from Gaussian distribution by reparameterizing $p_{w}$ as
@@
where $\Phi$ is the cumulative distribution function (CDF) of $\mathcal{N}(0,1)$ and $\Phi^{-1}$ is it’s inverse. As illustrated in Fig. 4, the each weight $w$ is then independently sampled as
@@
To represent positive and negative parameter values in the crossbar, a differential pair of PCM devices is used. Each differential PCM (DPCM) cell, whether in WP or NP has two devices, programmed to conductances $G^{+}$ and $G^{-}$. Let the $\sigma_{p}(G)$ and $\sigma_{r}(G)$ denote programming and read noise for a given device with conductance $G$. We denote the conductance of a cell in weight plane storing parameter $z_{w}$ as $G_{w}$, and the conductance of the noise plane cell as $G_{n}$.
Noise source: By programming the NP cells such that that $G_{n}^{+}\approx G_{n}^{-}\approx G_{n}$, the mean conductance read out from this cell almost vanishes, while the noise adds up. Programming noise has the dominant effect on the total noise of the noise plane cell as it is significantly larger than read noise as per Fig. 3. Accordingly, the generated noise parameters are be approximated as
@@
Figure 4: Top: Probability parameter $p_{w}$ as a function of $z_{w}$. $z_{w}$ is clipped to $|z_{clip}|=3$. The error in $p_{w}$ due to these clippings in negligible. Inset: Illustration of equivalence of Uniform sampling with $p_{w}$ and Gaussian sampling with $z_{w}$. Bottom: Detailed illustration of generating weight samples and MVM operation. Weight plane cells are read row-by-row. Simultaneously, a noise plane row is selected through stochastic arbitration (SA). Read pulse magnitudes of WP and NP are adjusted to restore WP parameters to their original scale. Their currents are summed as per Kirchoff’s law and binarized to complete the operation described in (12). The binarized value is used to determine the sign of inputs being accumulated, accomplishing the MAC operation.
@@
Mapping parameters to WP: We now discuss the mapping strategy of a WP cell, with weight $w$ and the parameter $z_{w}$. We first scale up $z_{w}$ by a factor $\kappa$ to get $G_{w}$ which allows the full use of the available conductance dynamic range of the device. It also reduces the impact of both programming and read noise in WP, ensuring that only the NP noise determines the generation of the weight samples [6].
@@ III-C Single Core Operations
Figure 5: Left: Detailed core design, expanding on Fig. 1 and 4 illustrating all the elements in the crossbar and the periphery. After accumulation, the products are transferred to the register, labeled here as ‘Reg’. This facilitates pipelining by enabling the subsequent set of matrix-vector multiplications (MVMs) to commence while the outputs from the previous MVMs are sequentially transported out of the core. Right: Current integration and voltage sensing circuit. Current is converted to voltage through charge integration with ping-pong mode capacitor. A current mirror-based current divider is used to reduce the integrating current to capacitor. Sense amplifier binarizes the voltage to generate the weight samples.
We now discuss the details of the inference operation of a single DPCM crossbar core (Fig. 5). The size of the crossbar in the core is $144\times 128$, with $128$ rows dedicated to WP and $16$ rows dedicated to NP. Each of these areas of the crossbar have a different mechanism for row-select, therefore they each have a dedicated decoder. A DPCM cell consists of two PCM devices each with its own transistor selector arranged as a differential pair, in the standard 1T1R configuration.
@@
The current from WP and NP is integrated as shown in Fig.  5. The current integrator, adopted from [44], consists of a current divider in the form of a mirror circuit, where the W/L ratio of the transistors involved can be adjusted to control the current ratio. This is done largely to reduce the magnitude of current integrating into the capacitor, enabling the use of smaller capacitors [13]. The differential sense amplifier then performs binarization based on the converted voltage, thereby implementing (16). As the capacitors need to be reset before being reused for integration, two capacitors in ping-pong fashion can be utilized. This way, while one capacitor discharges, the other can be employed for current integration. Upon integration, the sense amplifier, as described in [45] binarizes the converted voltage to generate $w\in\{-1,+1\}$.
@@
Multi-core operation: Since most typical layers in neural networks are larger than the size of a Bayes2IMC core ($144\times 128$), a layer is usually mapped across multiple cores [18]. We refer to a collection of such cores as a tile. Partial sums of such cores are then carried out across all the cores. The partial sums are then routed into a tile-level neuron processing unit performing pre-activation accumulation, BN, ReLU, and max-pooling.
@@ III-D Post-Processing Logit Correction
As described before, the non-idealities of the PCM devices lead to distorted logit distributions. We denote the RV for this distribution as $\tilde{L}_{k}=l(k|\boldsymbol{x},\tilde{\boldsymbol{w}})$, where $\tilde{\boldsymbol{w}}\in\mathcal{W}$ represents the hardware-sampled distorted model. The hardware distorted logit distribution also exhibits a bimodal structure made of distributions $\hat{p}(\tilde{L}_{k}|y=k)$ and $\hat{p}(\tilde{L}_{k}|y\neq k)$, which we model as Gaussian distributions with parameters $\tilde{\theta}_{k,1}=(\tilde{\mu}_{k,1},\tilde{\sigma}_{k,1})$ and $\tilde{\theta}_{k,0}=(\tilde{\mu}_{k,0},\tilde{\sigma}_{k,0})$, respectively.
The hardware distortions lead to reduction of resolvability of the predictions $p(y|\boldsymbol{w},w)$. This results in blurring of the modes, causing degradation of separability. Since the degree of corruption cannot be predicted in advance, we developed a post-facto algorithm that corrects these distortions after the network has been mapped to hardware, illustrated in Fig. 7. During the calibration phase, we use a subset $\mathcal{D}^{\mathrm{cal}}=\{x^{\mathrm{cal}}_{i},y^{\mathrm{cal}}_{i}\}_{i=1}^{N^{\mathrm{cal}}}$ of the available data, called calibration data, to run inference on the programmed network and estimate the Gaussian parameters $\tilde{\theta}_{k,1}$ and $\tilde{\theta}_{k,0}$.
We assume a one-to-one correspondence between $L_{y}$ and $\tilde{L}_{y}$ as an affine perturbation, such that both variables differ by scale and offset factors depending on parameters $\theta_{k,0}$, $\theta_{k,1}$, $\tilde{\theta}_{k,0}$, and $\tilde{\theta}_{k,1}$. If the true class $y$ is known during the inference in operational phase, then the correction can be described as
$\hat{l}_{k}=\mathbb{E}[L_{k}|\tilde{L}_{k},y]=\frac{(\tilde{L}_{k}-\tilde{\mu}_{k,c})}{\tilde{\sigma}_{k,c}}\sigma_{k,c}+\mu_{k,c},$ (17)
where $c=\delta_{\{y=k\}}$. However, since the true label $y$ is unknown at inference, we instead weight the correction by the estimated probability of $\{y=k\}$ (depicted in Fig. 7 panel (4)) using a Bayesian update as
@@
$\quad+\mathrm{Pr}(y\neq k|\tilde{L}_{k})\mathbb{E}[L_{k}|\tilde{L}_{k},y\neq k],$ (19)
where
$\mathrm{Pr}(y=k|\tilde{L}_{k})=\frac{\mathrm{Pr}(y=k)\mathrm{Pr}(\tilde{L}_{k}|y=k)}{\mathrm{Pr}(\tilde{L}_{k})}\\ =\frac{\mathrm{Pr}(y=k)\mathcal{N}(\tilde{L}_{k}|\tilde{\mu}_{k,1},\tilde{\sigma}_{k,1})}{\mathrm{Pr}(y=k)\mathcal{N}(\tilde{L}_{k}|\tilde{\mu}_{k,1},\tilde{\sigma}_{k,1})+\mathrm{Pr}(y\neq k)\mathcal{N}(\tilde{L}_{k}|\tilde{\mu}_{k,0},\tilde{\sigma}_{k,0})}$ (20)
and where $\mathrm{Pr}(y\neq k|\tilde{L}_{k})$ is computed in a similar fashion. For a balanced dataset, prior probabilities $\mathrm{Pr}(y=k)$ and $\mathrm{Pr}(y\neq k)$ are equal to $\frac{1}{n}$ and $\frac{n-1}{n}$, respectively. Accordingly, the posterior probabilities $\mathrm{Pr}(y=k|\tilde{L}_{k})$ and $\mathrm{Pr}(y\neq k|\tilde{L}_{k})$ determine the strength of correction application (scale and shift) towards either side, decreasing the fuzziness between the two modes. The process is repeated for each of the $N_{MC}$ predictors in the ensemble. As these operations are performed only on the output layer, they can be handled by a dedicated computational block off-tile, amortizing the overall cost.
@@ IV Experiments and Results
In this section, we evaluate the performance of the proposed Bayes2IMC architecture. We first train a VGGBinaryConnect network on CIFAR-10 dataset. VGGBinaryConnect is a modified version of a 9-layer frequentist VGG network with binary synapses described in [51] and [52]. Its BBNN counterpart was trained using the BayesBiNN optimizer described in [34]. For benchmarking, the frequentist version of the same network was trained using a Straight-Through Estimator (STE) with ADAM optimizer, as described in [34]. The BBNN network was then transferred to a custom PCM hardware simulator implementing the noise characteristics described in Fig. 3 as well as appropriate input and accumulator quantization. Inference was then performed with $N_{MC}=10$. The logit correction method discussed in sec. III-D were calibrated with $2000$ validation set images. All the results were reported from 6 different runs for each scenario, with each run comprising of (in case of a hardware implementation) an initial device programming step followed by inference. Accuracy, ECE, and uncertainty quantification results were obtained under various scenarios. Drift performance was evaluated for up to $T=10^{7}$ s. Finally, hardware projections were calculated for the Bayes2IMC architecture and compared with equivalent SRAM architecture and existing implementations in the literature.
@@ IV-A Accuracy and ECE Performance
Accuracy and ECE evaluated on hardware (with logits corrected and uncorrected) were benchmarked against FP32 software network, a frequentist STE-based network, and a fixed point (FxP) implementation (Fig. 9). The fixed point network was parametrized with $5$-bit parameters for a fair comparison with a PCM implementation, as the effective PCM bit capacity is reported to be $4$ [53] making DPCM capacity $5$-bits. The frequentist network was also transferred to PCM simulator for hardware simulation.
@@ IV-C Drift Mitigation
The effect of drift on accuracy and ECE was studied up to $T=10^{7}$ s, as shown in Fig. 11. We observe that accuracy as well as ECE degrades significantly due to conductance drift. With no correction, the net accuracy of the network comes down nearly to $10\%$ in $10^{6}$ s, which is equivalent to random guessing. In around $10^{5}$ s, i.e., slightly longer than a day, the accuracy reduces to $20\%$. Logit correction slightly improve the performance – with accuracy post correction at $10^{5}$ s being around $70\%$. It is worth noting that it takes around $10^{6}$ s for the accuracy of the output of Bayes2IMC to reduce to $10\%$ without any corrections. This is significantly longer than the time taken by analog PCM implementations such as [47], where the time taken for the accuracy to reduce to random guessing without any drift mitigation strategy is around $1000$ s. The reparametrization of $p_{w}$ to $z_{w}$ before mapping, rather than mapping $p_{w}$ directly, ensured significant robustness to drift, as seen in Fig. 8.
@@ IV-D Hardware Estimation
Total core area mm2 0.015 0.14
@@
Our simulations are based on large-array PCM hardware results reported in the literature by IBM Research where the memristive device was integrated with industry-standard 90 nm CMOS technology. Hence, we consider 90 nm node for our hardware projections incorporating the device noise characteristics based on PCM cells fabricated in this node [24]. The cell read power and area for PCM devices used here are based on [45]. SRAM read power is assumed to follow the trends reported in [54, 55], and scaled to equivalent numbers for the 90 nm node using scaling laws described in [56]. Cell area of a single PCM was taken to be $29$ F2 per device as per [45], whereas that of a 6T SRAM-based cell was taken to be $120$ F2 [18]. We consider a $128\times 128$ parameter storage per crossbar for both PCM and SRAM core. Additionally, $16$ rows are apportioned for NP, taking the total number of PCM core rows to $144$. While each row read in SRAM requires only a single clock cycle, the PCM architecture requires $8$ clock pulses for $n_{r}=1$ and $4$ clock pulses for $n_{r}=2$, as per Fig. 6. For the $n_{r}=2$ scenario, the read pulse width can shrink to $2$ clock pulses at around $10^{7}$s because of drift compensation.
@@
Fig. 13 shows the area breakdown of an PCM Bayes2IMC and the SRAM core respectively. In a Bayes2IMC PCM core, the digital peripheral circuitry—consisting of the accumulator, register, multiplexer block, and decoders—accounts for the majority of the overall area, making up $82\%$ of the footprint. In contrast, for SRAM, the crossbar area is comparable to the size of the peripheral area. Because of the dominance of peripherals, the total area of PCM core is only 2×2$\times$2 × smaller than that of SRAM even though the PCM crossbar is almost 14×14$\times$14 × smaller than its SRAM counterpart.
@@
The digital peripheral circuitry in latter also consists of the PRNG unit, which was designed with maximal reuse architecture discussed in [8], where all $4$-bytes on a $32$ bit LFSR were utilized for random number generation. The PRNG unit makes up almost $10\%$ of total peripheral area and $6\%$ of the total core area. Additionally, we highlight the area required to store $128$ pairs of BN coefficients, one per column.
Table II also compares the power consumed by the SRAM and the Bayes2IMC core. Owing to the increase in clock rates due to the decrease in $T_{NP}$, we have reported the power for three modes, indicated as $8/4/2$ corresponding to $n_{r}=1$, $n_{r}=2$, and for $T_{NP}/T_{WP}=2$ at $t=10^{7}$ for $n_{r}=2$ respectively.
@@
We also compare our work with other works on Bayesian neural network architectures reported in literature. These include implementations on CMOS (FPGA and ASIC), as well as on NVM devices such as DW-MTJ and SOT-MRAM utilizing nanoscale stochasticity. We compute the power efficiency in terms of GOPS/W and total efficiency in terms of GOPS/W/mm2. For fair comparison, we translate these numbers to 90 nm node equivalent using scaling laws described in [56]. We find that our design achieves one of the best accuracies on CIFAR-10 despite having single-bit weights. It also outperforms all CMOS architectures as well as the DW-MTJ (iso-node) implementation in terms of area and power efficiency. We also have 1.2$\times$ better best-case ($T_{NP}/T_{WP}=2$) and comparable nominal case (0.78$\times$ @ $T_{NP}/T_{WP}=4$) iso-node power efficiency compared to the recent state-of-the-art [9]. Note that the approach reported in[9] achieves high efficiency by performing ensembling only on selected layers, as well as by utilizing selector-less, high-speed, SOT-MRAM cells.

### v2

@@ III-A Overview
The parameters $\boldsymbol{p}_{\boldsymbol{w}}=\{p_{w}\}_{w\in\boldsymbol{w}}$ required for sample generation are stored in WP. The noise, provided by the NP, is due to the inherent device-to-device and cycle-to-cycle stochasticity represented by the state-dependent programming noise shown in Fig. 3. This Weight Plane-Noise Plane (WP-NP) crossbar array architecture introduced in [6], assigns $M$ rows for the weight probability parameters, and $L$ rows in the NP to generate the required stochasticity. We reuse the hardware resource for sampling so that $L\ll M$, and empirically show that this does not affect network accuracy, leading to improved area efficiency. Noise cells from NP and weight cells from WP combine to generate $w$. We now elaborate the details of this sampling utilizing device stochasticity.
@@ III-B Leveraging Device Stochasticity
Reparametrization: To generate network parameters as per (9), we require a source of Uniform noise. However, the naturally available stochasticity from the PCM-based NP in the array is Gaussian distributed.
Here, we develop a principled method for sampling binary weights from Gaussian distribution using transform sampling [41]. Accordingly, we reparametrize or transform the probability parameter $p_{w}$ into another parameter, $z_{w}$ such that it can be used with Gaussian noise for weight sampling. The reparametrization is performed as
@@
where $\Phi$ is the cumulative distribution function (CDF) of $\mathcal{N}(0,1)$ and $\Phi^{-1}$ is its inverse. As illustrated in Fig. 4, the each weight $w$ is then independently sampled as
@@
We use the differential pair configuration of PCM devices, a standard method for representing positive and negative parameters, as described in [23, 42, 43, 13]. Each differential PCM (DPCM) cell, whether in WP or NP has two devices, programmed to conductances $G^{+}$ and $G^{-}$. Let the $\sigma_{p}(G)$ and $\sigma_{r}(G)$ denote programming and read noise for a given device with conductance $G$. We denote the conductance of a cell in weight plane storing parameter $z_{w}$ as $G_{w}$, and the conductance of the noise plane cell as $G_{n}$.
Noise source: As opposed to the standard implementation of DPCM weights, here we programme the NP cells such that that the approximate equality $G_{n}^{+}\approx G_{n}^{-}\approx G_{n}$ holds. This ensures that the mean conductance read out from this cell almost vanishes, while the programming noise component from the devices adds up. Programming noise has the dominant effect on the total noise of the noise plane cell as it is significantly larger than read noise as per Fig. 3. Accordingly, the generated noise parameters are approximated as
@@
Figure 4: Top: Probability parameter $p_{w}$ as a function of $z_{w}$. $z_{w}$ is clipped to $|z_{clip}|=3$. The error in $p_{w}$ due to these clippings in negligible. Inset: Illustration of equivalence of Uniform sampling with $p_{w}$ and Gaussian sampling with $z_{w}$. Bottom: Detailed illustration of weight sampling as per (12), and MVM operation as per (16).
@@
Mapping parameters to WP: We now discuss the mapping strategy of a WP cell, with weight $w$ and the parameter $z_{w}$. We first scale up $z_{w}$ by a factor $\kappa$ to get $G_{w}$ which allows the full use of the available conductance dynamic range of the device. IMC also reduces the impact of both programming and read noise in WP, ensuring that only the NP noise determines the generation of the weight samples [6].
@@ III-C Single Core Operations
Figure 5: Left: A multi-tile core block diagram, with each tile composed of the Bayes2IMC cores along with peripherals such as BN unit, ReLU, pooling and buffers. Post-processing unit acts on the final layers for logit correction described in Sec.III-D.
Right: Detailed core design, expanding on Fig. 1 and 4 illustrating all the elements in the crossbar and the periphery.
We now discuss the details of the inference operation of a single DPCM crossbar core (Fig. 5). The size of the crossbar in the core is $144\times 128$, with $128$ rows dedicated to WP and $16$ rows dedicated to NP. The NP row size of $16$ was deemed to be sufficiently large to balance the need for reduced correlation amongst the parameters during sampling and area and power efficiency of the core [6]. Correlation was further reduced by utilizing stochastic arbitration, as discussed in Sec. III-B.
Each of these areas of the crossbar have a different mechanism for row-select, therefore they each have a dedicated decoder. A DPCM cell consists of two PCM devices each with its own transistor selector arranged as a differential pair, in the standard 1T1R configuration.
@@
Current integrator is then used to accumulate the current from WP and NP. A current integrator design adopted from [45] and [13] is depicted in Supplementary Section C. The source line currents are accumulated into integrating capacitors, one each for the two SLs corresponding to the differential pair. The output voltages of the capacitors thus developed are then binarized using a differential sense amplifier, thereby implementing (16).
@@
Multi-core operation: Since most typical layers in neural networks are larger than the size of a Bayes2IMC core ($144\times 128$), a layer is usually mapped across multiple cores [18]. We refer to a collection of such cores as a tile. Partial sums of such cores are then carried out across all the cores. The partial sums are then routed into a tile-level neuron processing unit performing pre-activation accumulation, Batch Normalization (BN), ReLU, and max-pooling.
@@ III-D Post-Processing Logit Correction
As described before, the non-idealities of the PCM devices lead to distorted logit distributions. The hardware distortions lead to reduction of resolvability of the predictions $p(y|\boldsymbol{w},\mathcal{D})$. This results in blurring of the modes, causing degradation of separability. The distorted distribution, with its random variable $\tilde{L}_{k}$, also exhibits a bimodal structure with Gaussian distributions $\hat{p}(\tilde{L}_{k}|y=k)$ and $\hat{p}(\tilde{L}_{k}|y\neq k)$. These distributions are parametrized by $\tilde{\theta}_{k,1}=(\tilde{\mu}_{k,1},\tilde{\sigma}_{k,1})$ and $\tilde{\theta}_{k,0}=(\tilde{\mu}_{k,0},\tilde{\sigma}_{k,0})$, respectively. Since the degree of corruption cannot be predicted in advance, we developed a post-facto algorithm that corrects these distortions after the network has been mapped to hardware, illustrated in Fig. 7.
During the calibration phase, we use a subset $\mathcal{D}^{\mathrm{cal}}=\{x^{\mathrm{cal}}_{i},y^{\mathrm{cal}}_{i}\}_{i=1}^{N^{\mathrm{cal}}}$ of the available data, called calibration data, to run inference on the programmed network and estimate the Gaussian parameters $\tilde{\theta}_{k,1}$ and $\tilde{\theta}_{k,0}$.
@@
We assume a one-to-one correspondence between $L_{y}$ and $\tilde{L}_{y}$ as an affine perturbation, such that both variables differ by scale and offset factors depending on parameters $\theta_{k,0}$, $\theta_{k,1}$, $\tilde{\theta}_{k,0}$, and $\tilde{\theta}_{k,1}$. Let us define the affine correction estimate for each mode $k$ as
$\mathbb{E}[L_{k}|\tilde{L}_{k},y]=\frac{(\tilde{L}_{k}-\tilde{\mu}_{k,c})}{\tilde{\sigma}_{k,c}}\sigma_{k,c}+\mu_{k,c},$ (17)
where $c=\delta_{\{y=k\}}$.
The correction is then applied by weighting the affine correction estimate by the estimated probability of $\{y=k\}$ (depicted in Fig. 7 panel (4)) using Bayesian update as
@@
$\quad+\mathrm{Pr}(y\neq k|\tilde{L}_{k})\mathbb{E}[L_{k}|\tilde{L}_{k},y\neq k].$ (19)
Here, the posteriors $\mathrm{Pr}(y=k|\tilde{L}_{k})$ and $\mathrm{Pr}(y\neq k|\tilde{L}_{k})$ are calculated using the Bayes’ theorem, with likelihood obtained from the value of Gaussian distribution function parametrized by $\tilde{\theta}_{k,1}$ and $\tilde{\theta}_{k,0}$ respectively at $\tilde{L}_{k}$.
For a balanced dataset, prior probabilities $\mathrm{Pr}(y=k)$ and $\mathrm{Pr}(y\neq k)$ are equal to $\frac{1}{n}$ and $\frac{n-1}{n}$, respectively. Further details are provided in Supplementary Section B. Accordingly, the posterior probabilities $\mathrm{Pr}(y=k|\tilde{L}_{k})$ and $\mathrm{Pr}(y\neq k|\tilde{L}_{k})$ determine the strength of correction application (scale and shift) towards either side, decreasing the fuzziness between the two modes. The process is repeated for each of the $N_{MC}$ predictors in the ensemble. As these operations are performed only on the output layer, they can be handled by a dedicated computational block off-tile, amortizing the overall cost.
@@ IV Experiments and Results
In this section, we evaluate the performance of the proposed Bayes2IMC architecture. We first train a VGGBinaryConnect network on CIFAR-10, following the standard practice in the literature for benchmarking Bayesian neural networks [10, 9, 50]. VGGBinaryConnect is a modified version of a 9-layer frequentist VGG network with binary synapses described in [51] and [52]. IMCs BBNN counterpart was trained using the BayesBiNN optimizer described in [34]. Adopting the hardware-aware training methodology introduced in [23], the network parameters are trained by injecting state-dependent Gaussian noise representative of PCM programming stochasticity, thus avoiding the need for expensive on-chip training. For benchmarking, the frequentist version of the same network was trained using a Straight-Through Estimator (STE) with ADAM optimizer, as described in [34]. The BBNN network was then transferred to a custom PCM hardware simulator implementing the noise characteristics described in Fig. 3 as well as appropriate input and accumulator quantization. Inference was then performed with $N_{MC}=10$. The logit correction method discussed in sec. III-D were calibrated with $2000$ validation set images. All the results were reported from 6 different runs for each scenario, with each run comprising of (in case of a hardware implementation) an initial device programming step followed by inference. Accuracy, ECE, and uncertainty quantification results were obtained under various scenarios. Drift performance was evaluated for up to $T=10^{7}$ s. Finally, hardware projections were calculated for the Bayes2IMC architecture and compared with equivalent SRAM architecture and existing implementations in the literature.
@@ IV-A Accuracy and ECE Performance
Accuracy and ECE evaluated on hardware (with logits corrected and uncorrected) were benchmarked against FP32 software BBNN (i.e. probability parameters are represented in FP32), a frequentist STE-based network, and a fixed point (FxP) implementation (Fig. 9). The fixed point network was parametrized with $5$-bit parameters for a fair comparison with a PCM implementation, as the effective PCM bit capacity is reported to be $4$ [53] making DPCM capacity $5$-bits. The frequentist network was also transferred to PCM simulator for hardware simulation.
@@ IV-C Drift Mitigation
The effect of drift on accuracy and ECE was studied up to $T=10^{7}$ s, as shown in Fig. 11. We observe that accuracy as well as ECE degrades significantly due to conductance drift. With no correction, the net accuracy of the network comes down nearly to $10\%$ in $10^{6}$ s, which is equivalent to random guessing. In around $10^{5}$ s, i.e., slightly longer than a day, the accuracy reduces to $20\%$. Logit correction slightly improve the performance – with accuracy post correction at $10^{5}$ s being around $70\%$. IMC is worth noting that it takes around $10^{6}$ s for the accuracy of the output of Bayes2IMC to reduce to $10\%$ without any corrections. The reparametrization of $p_{w}$ to $z_{w}$ before mapping, rather than mapping $p_{w}$ directly, ensured significant robustness to drift, as seen in Fig. 8.
@@ IV-D Hardware Estimation
Total crossbar area mm2 0.015 0.14
@@
Our simulations are based on large-array PCM hardware results reported in the literature by IBM Research where the memristive device was integrated with industry-standard 90 nm CMOS technology. Hence, we consider 90 nm node for our hardware projections incorporating the device noise characteristics based on PCM cells fabricated in this node [23]. The cell read power and area for PCM devices used here are based on [54]. SRAM read power is assumed to follow the trends reported in [55, 56], and scaled to equivalent numbers for the 90 nm node using scaling laws described in [57]. Cell area of a single PCM was taken to be $29$ F2 per device as per [54], whereas that of a 6T SRAM-based cell was taken to be $120$ F2 [18]. We consider a $128\times 128$ parameter storage per crossbar for both PCM and SRAM core. Additionally, $16$ rows are apportioned for NP, taking the total number of PCM core rows to $144$. While each row read in SRAM requires only a single clock cycle, the PCM architecture requires $8$ clock pulses for $n_{r}=1$ and $4$ clock pulses for $n_{r}=2$, as per Fig. 6. For the $n_{r}=2$, the read pulse width can shrink to $2$ clock pulses at $10^{7}$ s because of drift compensation.
@@
Fig. 13 shows the area breakdown of an PCM Bayes2IMC and the SRAM core respectively. In a Bayes2IMC PCM core, the digital peripheral circuitry—consisting of the accumulator, register, multiplexer block, and decoders—accounts for the majority of the overall area, making up $81\%$ of the footprint. In contrast, for SRAM, the crossbar area is comparable to the size of the peripheral area. Because of the dominance of peripherals, the total area of PCM core is only 2×2$\times$2 × smaller than that of SRAM even though the PCM crossbar is almost 10×10$\times$10 × smaller than its SRAM counterpart.
@@
The digital peripheral circuitry in latter also consists of the PRNG unit, which was designed with maximal reuse architecture discussed in [8], where all $4$-bytes on a $32$ bit LFSR were utilized for random number generation. The PRNG unit makes up almost $10\%$ of total peripheral area and $6\%$ of the total core area. Additionally, we highlight the area required to store $128$ pairs of BN coefficients, one per column. Table II also compares the power consumed by the SRAM and the Bayes2IMC core. Owing to the increase in clock rates due to the decrease in $T_{NP}$, we have reported the power for three modes, indicated as $8/4/2$ corresponding to $n_{r}=1$, $n_{r}=2$, and for $T_{NP}/T_{WP}=2$ at $t=10^{7}$ for $n_{r}=2$ respectively.
@@
We also compare our work with other works on Bayesian neural network architectures reported in literature. These include implementations on CMOS (FPGA and ASIC), as well as on NVM devices such as DW-MTJ and SOT-MRAM utilizing nanoscale stochasticity. We compute the power efficiency in terms of GOPS/W and total efficiency in terms of GOPS/W/mm2. For fair comparison, we translate these numbers to 90 nm node equivalent using scaling laws described in [57]. We find that our design achieves one of the best accuracies on CIFAR-10 despite having single-bit weights. IMC also outperforms all CMOS architectures as well as the DW-MTJ (iso-node) implementation in terms of area and power efficiency. We also have 1.2$\times$ better best-case ($T_{NP}/T_{WP}=2$) and comparable nominal case (0.78$\times$ @ $T_{NP}/T_{WP}=4$) iso-node power efficiency compared to the recent state-of-the-art [9]. Note that the approach reported in[9] achieves high efficiency by performing ensembling only on selected layers, as well as by utilizing selector-less, high-speed, SOT-MRAM cells.